How are addition and multiplication use to solve a linear equation? Demonstrate by solving "(1/5) x - (2/3) = 6x(2/5)"

Answer by mananth(16946)

(1/5) x - (2/3) = 6x(2/5)
..
Add +2/3 to both sides
(1/5) x - (2/3)+2/3 = 6x(2/5)+2/3
..
(1/5) x = 6x(2/5)+(2/3)
..
add - 6x(2/5) on both sides
(1/5)x -6x(2/5) =6x(2/5) -6x(2/5)+(2/3)
..
(1/5)x -6x(2/5) = 2/3
LCM = 5 on left side
x-12x /5 =2/3
-11x/5 =2/5
multiply by 5
-11x/5 *5 = 2/5 *5
-11x=2
divide by -11
-5x/-11= 2/-11
x= - 2/11